OpenStudy (anonymous):

Choose the correct GCF for the following trinomial: 12x2 − 18x + 12.

OpenStudy (anonymous):

6?

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

12=6*2 18=6*3 12=6*2 thus 6

OpenStudy (anonymous):

Ty